OpenStudy (anonymous):

what is the square root of 72x^2

OpenStudy (anonymous):



OpenStudy (anonymous):

\[6*\sqrt{2*}x\]

OpenStudy (anonymous):

\[6\sqrt{2}x\]

OpenStudy (hoblos):

√72x^2 = x√(36*2) = 6x√2

OpenStudy (anonymous):

 What about this one?